Two planes leave Honolulu International Airport at the same time. One heads towards the Big Island at 195 mph. the other heads in the opposite direction towards Kauai at 220 mph. How far apart are they after 20 minutes?
Answer by ptaylor(2198)

Distance(d)=rate(r) times time(t) or d=rt
Clearly, the planes are separating at the rate of 220 mph plus 195 mph or
415 mi/hr. Also 20 min =1/3 hr----So:
d=rt
d=415 mi/hr(1/3 hr)
d=138.333 mi
Another approach would be to determine how far each plane traveled in 20 min and then add the two distances together:.e.g.,
195/3+220/3=
65+73.333=138.333 mi
Hope this helps----ptaylor